# Automatic Hardware Pragma Insertion in High-Level Synthesis: A Non-Linear Programming Approach

### v2

@@ 2.4. Overview of NLP-DSE
Although NLP-DSE it can be used to compute pragmas without any DSE, the inherent limit of analytical models persist with NLP-DSE : as the implementation details of the back-end toolchain for HLS and synthesis may not be accurately captured by a model, DSE remains needed for best performance.
NLP-DSE enables the exploration of different parallelism and configuration spaces by constraining the level of parallelism, as detailed in Section 6. Our model is presented in Sec. 3, and we prove it is a performance lower bound in Sec. 4, an important feature to be able to prune designs during the search without the risk of losing performance. The associated NLP formulation is provided in Section 5.
The effectiveness of our framework is demonstrated in Table 3, comparing the performance and time-to-solution of NLP-DSE . We also display the result of the first synthesizable design produced, NLP-DSE -FS.
@@
Table 3. Comparison of NLP-DSE , NLP-DSE -FS, which provides the result of the first synthesizable design, the source-to-source compiler Merlin without pragma insertion, and AutoDSE in terms of throughput (GF/s), DSE time (mn) and DSP utilization (%) for the kernels 2mm, Gemm, and Gramschmidt.
@@
For 2mm, the first design synthetizable allows us to have a better QoR vs. AutoDSE but our DSE shows its interest in finding a configuration 8.9 times faster than the first configuration found by the DSE. More detail can be found in Section 8.
It is noteworthy that our Design Space Exploration (DSE) approach, detailed in Section 6, deviates intentionally from AutoDSE. Unlike AutoDSE, which starts with a pragma-free configuration and gradually introduces pragmas, we begin with configurations characterized by the lowest theoretical latency, emphasizing high levels of parallelism. This deliberate departure from the conventional approach is further discussed in Section 6.
In Section 7, we present a comprehensive evaluation of our framework, demonstrating the improvements over AutoDSE that can be achieved by specializing to affine programs, in terms of design throughput and time-to-solutions across various benchmarks. The results indicate an average performance improvement of 5.69x and 17.24x in terms of DSE time and design throughput, respectively, with only a marginal decrease in throughput for 1 out of 47 benchmarks. Importantly, the time-to-solution of NLP-DSE it is consistently up to 30x faster than AutoDSE across all benchmarks.
@@ 3. Modeling Programs and Their Pragmas
We now present our analytical performance model. We assume the input programs are polyhedral programs (Feautrier, 1992; Girbal et al., 2006), and therefore exact loop trip counts can be computed by static analysis, similarly for all data dependencies.
@@ 3.1. Program Representation
We employ the code below as a running example with the pragma above the loops as AMD/Xilinx Merlin. For presentation simplicity, we assume each loop iterator in the program region has been renamed to a unique name, so that we can uniquely identify loops in the code by their iterator name.
@@ 9 S3(i,j2);
The summary AST for this loop is simply built by creating one node per for loop and one per statement Sx, the body of a loop is made of loops and/or statements, and their nodes are children of said loop in the tree, listed in their syntactic order. For example, above, utilizing a constructor notation, it gives: $Loop_{i}(Loop_{j1}(S1),Loop_{j2}(S2,S3))$. Then, a simple rewrites of this tree using loop properties and composition operators will lead to the proposed analytical model, as outlined below. We first describe the loop properties we associate to each loop. We consider combinations of the following pragmas, based on Merlin’s optimizations, for the loop with iterator i:
@@
We therefore associate to each loop i in the program a property vector that informs about the optimizations to be considered. We define $\vec{P}V_{i}$ as follows: $\vec{P}V_{i}:<ispipelined_{i},II_{i},uf_{i},tile_{i},TC^{min}_{i},TC^{max}_{i}>$ where we have: $ispipelined_{i}=1$ if the loop is pipelined, 0 otherwise; $II_{i}$ is the initiation interval, set to 1 by default; $uf_{i}$ is the parallelism/unroll factor, set to 1 by default (no #pragma ACCEL parallel pragma) and set to $TC^{max}_{i}$ if parallel is defined without a factor $uf_{i}$ specified. $tile_{i}$ is the TC of the innermost loop after strip mining. $TC^{min}_{i}$ is the minimal trip count of loop i, for any of its execution in the program. We also compute the maximal trip count over all executions. These values are computed using polyhedral analysis on the loops (Pouchet et al., 2013b). The pragma cache transfers above the loop i the data needed for the computation of this loop nest for the array a.
@@
Modeling Vitis optimizations
AMD/Xilinx Vitis will apply several optimizations automatically, such as auto-pipeline and auto-loop-flatten, some other optimizations when the user gives a compilation option such as tree reduction.
Only a loop with a constant $TC$, i.e., $TC_{max}=TC_{min}$ can be unrolled. The unroll pragma options allow to specify the unrolling factor, $uf_{i}$. When the factor is not specified, it implies that the factor is equal to the $TC$ of the loop. When a loop is pipelined, all innermost loops are automatically fully unrolled. Hence, we also propagate unrolling information, e.g., to mark a full loop nest for full unrolling if an outer loop is marked with #pragma ACCEL pipeline. The pipeline pragma options allow to specify the objective II the user wants to achieve. When II is not specified, it is automatically set to 1. In addition, Vitis will auto-pipeline with a target II of 1 the innermost loops which are not fully unrolled for each nested loop.
Within Vitis, users can enable optimizations like logarithmic time reduction through tree reduction. This optimization choice will be a global option within our model, applicable across the entire model rather than being limited to a specific loop.
Modeling Merlin optimizations
Finally, Merlin will also add automatic optimizations. It will explicitly strip-mine a loop when it is partially unrolled with the innermost loop having a $TC$ equal to the factor and the unrolled pragma applied to that loop. Similarly to Vitis, Merlin will auto-pipeline. Merlin also applies a program transformation for certain pragmas. When there are two perfectly nested and partially unrolled loops, Merlin swaps the two loops strip-mined (if legal), unrolled innermost and flatten and pipeline the two outermost loops. Further, Merlin will automatically transfer the data from off-chip to on-chip and cache on-chip with packing, with a maximum packing of 512 bits for our FPGA while computing if the footprint of the data fit on-chip by static analysis. The pragma tile allows Merlin to strip mine a loop and give the compiler the opportunity to transfer less/more data while respecting resource constraints. For our model we suppose an optimistic data transfer i.e., all memory transfers are done with a packing of 512 bits and each data are transferred once (perfect data reuse).
@@
4. Modeling of Latency and Resource Lower Bound
We present the foundational components of our analytical performance model and demonstrate how this model computes a lower bound on latency while adhering to resource constraints. The theorems and their accompanying proofs can be found in Appendix B.
4.1. Analytical Model Template
Consider the expression $I^{\vec{P}V_{i}}_{i}(X)$ for the loop $i$, where $X$ denotes a subpart of the program, and the operators are defined as follows:
$I^{\vec{PV}_{i}}_{i}(X)=\begin{cases}\left\lfloor II_{i}\cdot\left(\frac{TC^{avg}_{i}}{uf_{i}}-\text{ispip}_{i}\right)\right\rfloor+X&\text{if loop }i\text{ is pipelined}\\ &(\text{ispip}_{i}=1)\\ \\ \left\lfloor II_{i}\cdot\left(\frac{TC^{avg}_{i}}{uf_{i}}-\text{ispip}_{i}\right)\right\rfloor\cdot X&\text{otherwise }(\text{ispip}_{i}=0)\end{cases}$
For simplicity, we will represent the operation as $\odot$ to differentiate between the two different cases.
Let $C_{i}^{\vec{P}V_{i}}(X_{1},X_{2},\dots,X_{n})$ which compose the different sub part of a program under the loop $i$.
$C_{i}^{\vec{P}V_{i}}(X_{1},X_{2},\dots,X_{n})=\begin{cases}max(X_{1},X_{2},\dots,X_{n})&\text{ if }(X_{1},\dots,X_{n})\text{ do not have dependencies (WaR, RaW, WaW)}\\ \sum_{k=1}^{n}X_{k}&\text{otherwise}\end{cases}$
1S0: x1 += y;
2S1: x2 += x1 + z;
Listing 2: Example of code where two statements have dependencies
1S0: x1 += y;
2S1: x2 += z;
Listing 3: Example of code where two statements do not have dependency
In List 2, statements $S0$ and $S1$ exhibit dependencies, resulting in a program latency equal to the sum of their latencies, corresponding to $issequence=1$. Conversely, the statements in List 3 lack dependencies, allowing $S0$ and $S1$ to be executed concurrently. Therefore, the latency is determined by the maximum of $S0$ and $S1$ latencies, representing the case $inparallel=1$.
Finally, $SL^{\vec{P}V_{i}}_{i}(\vec{S}_{k})$ denotes a region of straight-line code (e.g., an inner-loop body). Intuitively, $SL$ will represent a lower bound on the latency of a code block such that by composition across all loops as per the template formula, the result remains a lower bound on the full program latency.
Given the summary AST of the program in constructor form, and the set of $\vec{P}V_{i}$ vectors for each loop, we build the analytical formula template as follows:
  1. (1)
    replace the $Loop_{i}(\vec{X})$ operator by $I^{\vec{P}V_{i}}_{i}(\vec{X})$ operator;
  2. (2)
    replace the $\vec{X}$ list by $C^{\vec{P}V_{i}}_{i}(\vec{X})$;
  3. (3)
    replace statement lists inside loop i $\vec{S}_{x}$ by $SL^{\vec{P}V_{i}}_{i}(\vec{S}_{x})$.
For our example, we can rewrite the expression:
$Loop_{i}(Loop_{j1}(S1),Loop_{j2}(S2,S3))$
as
$I^{\vec{P}V_{i}}_{i}(C^{\vec{P}V_{i}}(I^{\vec{P}V_{j1}}_{j1}(SL^{\vec{P}V_{j1}}_{j1}(S1)),I^{\vec{P}V_{j2}}_{j2}(SL^{\vec{P}V_{j2}}_{j2}(S2,S3))))$
Upon substituting the definitions and assuming that the statements are independent, we obtain:
$\left\lfloor II_{i}\cdot\left(\frac{TC^{avg}_{i}}{uf_{i}}-\text{ispip}_{i}\right)\right\rfloor\odot\max\left(\left\lfloor II_{j1}\cdot\left(\frac{TC^{avg}_{j1}}{uf_{j1}}-\text{ispip}_{j1}\right)\right\rfloor\odot SL^{\vec{P}V_{j1}}_{j1}(S1),\left\lfloor II_{j2}\cdot\left(\frac{TC^{avg}_{j2}}{uf_{j2}}-\text{ispip}_{j2}\right)\right\rfloor\odot SL^{\vec{P}V_{j2}}_{j2}(S2,S3)\right)$

### v4

@@ 3. Overview of NLP-DSE
Although NLP-DSE can be used to compute pragmas without any DSE, the inherent limits of analytical models persist with NLP-DSE: as the implementation details of the back-end toolchain for HLS and synthesis may not be accurately captured by a model. Hence, DSE remains needed for best performance. NLP-DSE enables the exploration of different parallelism and configuration spaces by constraining the level of parallelism, as detailed in Section 7. Our model is presented in Section 4, and we prove it is a performance lower bound in Section 6, an important feature to be able to prune designs during the search without the risk of losing performance. The associated NLP formulation is provided in Section 5.
The effectiveness of our framework is demonstrated in Table 4, which compares the performance and time-to-solution of NLP-DSE. Additionally, we present the results of the first synthesizable design produced, referred to as NLP-DSE-FS.
@@
Imp. vs. HARP 1x 13x 0.96x 2.74x N/A N/A
Table 4. Comparison of NLP-DSE, NLP-DSE-FS (which provides the result of the first synthesizable design), the source-to-source compiler Merlin without pragma insertion, AutoDSE and HARP, in terms of throughput (GF/s), DSE time (minutes), and DSP utilization (%) for the 2mm, Gemm, and Gramschmidt kernels using float and double data types
@@
For 2mm, the first design synthetizable allows us to have a better QoR vs. AutoDSE but our DSE demonstrates the ability to find a configuration 8.9 times faster than the first configuration found by the DSE. More detail can be found in Section 9.
It is noteworthy that our Design Space Exploration (DSE) approach, detailed in Section 7, deviates significantly from AutoDSE. Unlike AutoDSE, which starts with a pragma-free configuration and gradually introduces pragmas, we begin with configurations characterized by the lowest theoretical latency, emphasizing high levels of parallelism. This deliberate departure from the conventional approach is further discussed in Section 7.
In Section 8, we present a comprehensive evaluation of our framework, demonstrating the improvements over AutoDSE that can be achieved by specializing to affine programs, in terms of design throughput and time-to-solutions across various benchmarks. The results indicate an average performance improvement of 5.69x and 17.24x in terms of DSE time and design throughput, respectively, with only a marginal decrease in throughput for 1 out of 47 benchmarks. Importantly, the time-to-solution of NLP-DSE it is consistently up to 30x faster than AutoDSE across all benchmarks evaluated.
@@ 4. Modeling Programs and Their Pragmas
We now present our analytical performance model. We assume the input programs are affine programs (Feautrier, 1992; Girbal et al., 2006), and therefore exact loop trip counts can be computed by static analysis, similarly for all data dependencies.
@@ 4.1. Program Representation
We employ the code below as a running example with the pragma above the loops as AMD/Xilinx Merlin. For presentation simplicity, we assume each loop iterator in the program region has been renamed to a unique name, so that we can uniquely identify loops by their iterator name.
@@ 9 S3(i,j2);
Listing 2: Running example with the pragma above the loops as AMD/Xilinx Merlin
The summary AST for this program is simply built by creating one node per for loop and one per statement Sx, the body of a loop is made of loops and/or statements, and their nodes are children of said loop in the tree, listed in their syntactic order. For the example above, using a tree constructor notation, we get: $Loop_{i}(Loop_{j1}(S1),Loop_{j2}(S2,S3))$. Then, a simple rewrite of this formula by substituting the loop and statements by carefully chosen composition operators and descriptors of loop properties will lead to our non-linear analytical model, as outlined in Sec. 6.
We first describe the loop properties we associate to each loop. We consider combinations of the following pragmas, based on Merlin’s optimizations, for the loop with iterator i:
@@
We therefore associate to each loop i in the program a property vector that informs about the optimizations to be considered. We define $\vec{P}V_{i}$ as follows: $\vec{P}V_{i}:<ispipelined_{i},II_{i},uf_{i},tile_{i},TC^{min}_{i},TC^{max}_{i}>$ where we have: $ispipelined_{i}=1$ if the loop is pipelined, 0 otherwise; $II_{i}$ is the initiation interval, set to 1 by default; $uf_{i}$ is the parallelism/unroll factor, set to 1 by default (no #pragma ACCEL parallel pragma) and set to $TC^{max}_{i}$ if parallel is defined without a factor $uf_{i}$ specified. $tile_{i}$ is the TC of the innermost loop after strip mining. $TC^{min}_{i}$ is the minimal trip count of loop i, for any of its execution in the program. We also compute the maximal trip count over all executions. These values are computed using polyhedral analysis on the loops (Pouchet et al., 2013). The pragma cache transfers above the loop i the data needed for the computation of this loop nest for the array a.
@@
Modeling Vitis optimizations  AMD/Xilinx Vitis will apply several optimizations automatically, such as auto-pipeline and auto-loop-flatten, some other optimizations when the user gives a compilation option such as tree reduction. Only a loop with a constant trip count ($TC$), i.e., $TC_{max}=TC_{min}$ can be unrolled. The unroll pragma options allow to specify the unrolling factor, $uf_{i}$. When the factor is not specified, it implies that the factor is equal to the $TC$ of the loop. When a loop is pipelined, all innermost loops are automatically fully unrolled. Hence, we also propagate unrolling information, e.g., to mark a full loop nest for full unrolling if an outer loop is marked with #pragma ACCEL pipeline. The pipeline pragma options allow to specify the objective II the user wants to achieve. When II is not specified, it is automatically set to 1. In addition, Vitis will auto-pipeline with a target II of 1 the innermost loops which are not fully unrolled for each nested loop. Within Vitis, users can enable optimizations like logarithmic time reduction through tree reduction. This optimization choice will be a global option within our model, applicable across the entire model rather than being limited to a specific loop.
Modeling Merlin optimizations  Merlin applies automatic optimizations to enhance performance. When a parallel pragma is used, the loop undergoes explicit strip-mining during partial unrolling. The innermost loop’s trip count matches the unrolling factor, with the parallel pragma applied at this level. Similar to Vitis, Merlin automatically pipelines loops. Merlin also performs program transformations for specific pragmas. For example, in the case of two perfectly nested and partially unrolled loops, Merlin (when legal) swaps the strip-mined loops, unrolls the innermost loop, and flattens and pipelines the two outermost loops. Additionally, Merlin automates data transfers from off-chip to on-chip memory, caching data on-chip with packing. Our FPGA supports a maximum packing size of 512 bits. Merlin uses static analysis to ensure the data footprint fits on-chip. The tile pragma enables loop strip-mining, allowing the compiler to transfer more or less data while adhering to resource constraints. For our model, we assume an optimistic data transfer scenario—memory transfers occur with 512-bit packing, and each piece of data is transferred only once, reflecting perfect data reuse.
@@
5. Non-Linear Formulation for Pragma Insertion
Let $\mathcal{P}$ be the set of different possible pipeline configurations. Let $\forall p\in\mathcal{P}$ define $\mathcal{L}^{p}_{pip}$ the set of loops pipelined and $\forall l\in\mathcal{L}^{p}_{pip}$, $\mathcal{L}^{p}_{under\_pip_{l}}$ the set of loops within a loop pipelined and $\mathcal{L}^{p}_{above\_pip_{l}}$ the set of loops above the loop pipelined $l$. Let $\forall s\in\mathcal{S}$ define the set of nested loops which iterate the statement $s$, $\mathcal{L}_{s}$. $\forall a\in\mathcal{A}$ and for $d$ a dimension of the array $a$, let $\mathcal{C}_{a_{d}}$ be the set of loops which iterates the array $a$ at the dimension $d$. $\forall l\in\mathcal{L}$, let designate $d_{l}$ the maximum dependency distance of the loop $l$. And let $II_{s}$ be the II of the loop pipelined for the statement $s$.
The II for each loop, the dependencies, the properties of the loops, the trip count (TC), the iteration latency of the parallel operations and the reduction operations and the number of DSPs per operation per statements are computed at compile time with PolyOpt-HLS (Pouchet et al., 2013) and used as constants in the NLP problem.
Table 5, summarize the sets, variable and constants we use.
5.1. Variables
(1) $\forall l\in\mathcal{L},1\leq loop_{l}\_UF\leq TC_{l}$
(2) $\forall l\in\mathcal{L},1\leq loop_{l}\_tile\leq TC_{l}$
(3) $\forall l\in\mathcal{L},loop_{l}\_pip\in\{0,1\}$
Set Description
$\mathcal{L}$ the set of loops
$\mathcal{A}$ the set of arrays
$\mathcal{S}$ the set of statements
$\mathcal{O}_{s}$ the list of operations of the statement $s$
$\mathcal{O}_{s_{par}}$ the operations which can be done in parallel, i.e., do not have any loop-carried dependence
$\mathcal{P}$ the set of different possible pipeline configurations
$\mathcal{L}^{p}_{pip}$ the set of pipelined loops
$\mathcal{L}^{p}_{under\_pip_{l}}$ the set of loops under a pipelined loop
$\mathcal{L}^{p}_{above\_pip_{l}}$ the set of loops above the pipelined loop $l$
$\mathcal{L}_{s}$ the set of nested loops which iterate over the statement $s$
$\mathcal{C}_{a_{d}}$ the set of loops which iterate over the array $a$ at dimension $d$
$d_{l}$ the maximum dependency distance of the loop $l$
$AP_{a,d}$ Array Partition for the array $a$ in dimension $d$
Variable Description
$loop_{l}\_UF$ Unroll factor of the loop $l$
$loop_{l}\_tile$ Tile size of the innermost loop after strip-mining the loop $l$
$loop_{l}\_pip$ Boolean to indicate if the loop $l$ is pipelined
$loop_{l}\_cache\_array\_a$ Boolean to indicate if the array $a$ is transferred on-chip before the loop $l$
Constant Description
$TC_{l}$ Trip Count of the loop $l$
$II_{l}$ Initiation Interval (II) of the loop $l$
$IL_{par}$ Iteration Latency of the operations without dependencies in the statement $s$
$IL_{red}$ Iteration Latency of the operations with dependencies in the statement $s$
$DSP_{s_{op}}$ Number of DSPs used for the statement $s$ for the operation $op$
$DSP_{available}$ Number of DSPs available for the FPGA used
$max_{part}$ Maximum array partitioning defined by the user or the DSE (cf. Section 7)
$D_{a}$ Number (Integer) of dimensions of the array $a$
5.2. Modeling Compiler Transformations
Pipeline Rules:  Vitis HLS unrolls all loops under the pipelined loop. This implies that all loop $l$ under the pipelined loop must have $loop_{l}\_UF==TC_{l}$ (Eq. 5). Considering this constraint, it is important to note that for each statement, only one of the loops that iterate the statement can have a pragma pipeline (Eq. 6). Thanks to PolyOpt-HLS (Pouchet et al., 2013), we have the schedule of the kernel, including the loop order and which loops iterate over other loops. If multiple pipelined loops were present, the loops beneath the first pipelined loop would be unrolled instead. For example, in Listing 1, if Loop4 (line 8) is pipelined, this implies that Loop5 is fully unrolled with an unroll factor of 190.
(5) $\forall p\in\mathcal{P},\forall lp\in\mathcal{L}^{p}_{pip},\forall lbp\in\mathcal{L}_{under\_pip_{lp}},loop_{lp}\times loop_{lbp}\_UF==loop_{lp}\times TC_{lbp}$
Memory Transfer Rules:  Merlin automates the process of transferring data on-chip and applying array partitioning. The tool caches data on-chip and packs it in chunks of up to 512 bits, enabling efficient transfer speed. When the data is already present on-chip it can be reused provided that resource constraints are satisfied. The compiler caches on-chip the data only above the loop pipelined (Eq. 7).
If Loop4 in Listing 1 is pipelined, we need to cache at least all the data from the second dimension of D and tmp (i.e., the tiles $1\times 220$ and $1\times 190$) and all the elements of C (i.e., the tile $190\times 220$) on board to ensure efficient pipelining and unrolling.
(7) $\forall p\in\mathcal{P},\forall lp\in\mathcal{L}^{p}_{pip},\forall lbp\in\mathcal{L}_{under\_pip_{lp}}\forall a\in\mathcal{A},loop_{lbp}\_cache\_array_{a}==0$
Dependencies:  Loop-carried dependencies are managed using constraints (Eq 8). If a loop has a dependency distance of $n$, this means that if we unroll with an unroll factor $uf>n$ this is equivalent to unrolling the loop with a factor $uf=n$ because the statements corresponding to the iteration $\{uf+1,\dots,n\}$ will be executed only after the first n statements are executed due to the dependency. Loop-independent data dependencies are managed at the objective function, as elaborated in Section 5.3.
If we encounter code like $for(j=2;j<N;j++)y[j]=y[j-2]+3;$, a straightforward approach to handling this type of dependency is to impose a constraint such as $loop_{l}\_UF\leq 2$, which is represented by equation 8. In this case, due to dependencies an $UF>2$ is similar to $UF=2$.
(8) $\forall l\in\mathcal{L},\text{ if }dd_{l}>1,loop_{l}\_UF\leq d_{l}$
Array Partitioning:  The array partitioning is constrained (Eq. 9) to adhere to the maximum array partitioning limit (e.g., 1024 for AMD/Xilinx). Equations 10 and 11 enforce that the array partitioning must be equal to or a multiple of the unroll factor for dimension $d$. Equation 9 constrains the total array partitioning. This constraint limits the maximum unrolling factor while offering more flexibility than directly restricting individual unroll factors. By constraining total array partitioning, it indirectly controls the unroll factor by limiting the product of unroll factors, rather than imposing direct limits on each one.
For example, if we set $max_{part}$ to 1024, the array partitioning of array $D$ in Listing 1 will limit the product of the unroll factors of Loop3 and Loop4 to 1024. Specifically, $AP_{D,0}=\text{loop}_{Loop3}\_UF$ and $AP_{D,1}=\text{loop}_{Loop4}\_UF$, and the product across dimensions $d\in\llbracket 1,D_{a}\llbracket$ is constrained by $\prod_{d\in\llbracket 1,D_{a}\llbracket}AP_{a,d}=\text{loop}_{Loop3}\_UF\times\text{loop}_{Loop4}\_UF\leq max_{part}$.
(9) $\forall a\in\mathcal{A},\prod_{d\in\llbracket 1,D_{a}\llbracket}AP_{a,d}\leq max_{part}$
(10) $\forall a\in\mathcal{A},\forall d\in\llbracket 1,D_{a}\llbracket,\forall l\in C_{a_{d}},AP_{a,d}\%loop_{l}\_UF==0$
(11) $\forall a\in\mathcal{A},\forall d\in\llbracket 1,D_{a}\llbracket,\forall l\in C_{a_{d}},AP_{a,d}\geq loop_{l}\_UF$
Supplementary Rules:  In addition, we add the constraints for the divisibility of the problem size of the unroll factors (Eq. 12) and the tile size (Eq. 13). For this last, this corresponds to adding an upper bound to the product of the unroll factors of all the loops which iterate the same array on different dimensions.
During our DSE, we can force the solution to be fine-grained. In this case we add a constraint where the loop above the loop pipelined has a UF of 1, i.e., for all loop $l$ above the pipelined loop $loop_{l}\_UF==1$ (Eq. 14).
(12) $\forall l\in\mathcal{L},loop_{l}\_UF\%TC_{l}==0$
(13) $\forall l\in\mathcal{L},loop_{l}\_tile\%TC_{l}==0$
(14) $(Opt.)\forall l\in\mathcal{L},\forall l^{\prime}\in\mathcal{L}_{above\_l},loop_{l}\_pip*loop_{l^{\prime}}\_UF<=1$
We also constrain the resources, modeling their sharing optimistically. We consider the number of DSPs (Eq. 15) and on-chip memory (Eq. 16) used. As the consumption of DSPs can be difficult to estimate due to resource sharing we utilize an optimistic estimate, which considers a perfect reuse/sharing: as soon as a computation unit is free, its resource can be reused.
(16) $\sum_{a\in\mathcal{A}}\sum_{l\in\mathcal{L}}loop_{l}\_cache\_array_{a}\times footprint\_array_{a}\_loop_{l}\leq Mem$
5.3. Objective function
In every loop nest, there will invariably be a pipeline loop due to either user-inserted or compiler-inserted instructions (AMD/Xilinx Merlin and Vitis automatically insert the pragma pipeline if it is not done by the user or the previous compiler). Therefore, the objective function takes the following form: $TC_{ap}\times(IL+II\times(\frac{TC}{UF}-1))$, where $TC_{ap}$ includes the loops situated above the pipeline. Parallel loops above the pipeline can be coarse-grained parallelized. The iteration latency within the unrolled loop body is divided into either reduction operations or non-reduction operations, as reduction operations require logarithmic time for the reduction process. The variable IL encompasses the latencies of the statements found within the pipelined loop body. Independent statements can be executed in parallel and statements with dependencies are summed.
5.4. Example
1Loop0: for(i=0; i<2100; i++)
2 S0:y[i] = 0;
3Loop1: for(i=0; i<1900; i++) {
4 S1:t[i] = 0;
5 Loop2: for(j=0; j<2100; j++)
6 S2:t[i]+=A[i][j]*x[j];
7 Loop3: for(j=0; j<2100; j++)
8 S3:y[j]+=A[i][j]*t[i];
9}
Listing 3: AtAx code for Large problem size: $t=A*x;y=A^{t}*t$
We now use the AtAx kernel (Listing 3) as an illustration.
$S0$ and $S3$ do not have inter-iteration dependencies within their respective loops, $Loop0$ and $Loop3$. Therefore, it is possible to pipeline $Loop0$ and $Loop3$ with an initiation interval (II) $\geq 1$. In other words, there are no dependencies on previous iterations within the same loop for all iterations in these loops.
Loop1 and Loop2 are reduction loops, where the reduction operation is an addition with a latency of $IL_{+}$ cycles. Thus, the II for these loops must be greater or equal than $IL_{+}$.
If Loop1 is pipelined, there is a dependency between $S1$, $S2$, and $S3$. The dependency distance between $S1$ and $S2$ is 1, so we simply add $IL_{S1}$ and $IL_{S2}$. Between $S2$ and $S3$, the array $t$ can be read in statement $S3$ after the reduction of Loop2 is completed in $\log(N)$ cycles due to the tree reduction. Therefore, the final equation for the loop body cycle will be $IL_{S1}+IL_{S2}\times\log(N)+IL_{S3}$.
If statements can be executed simultaneously (i.e., there is no dependency), we use the maximum instead of an addition.
6. Latency and Resource Lower Bound
We now revisit the analytical modeling of programs in NLP-DSE, demonstrating that the objective function presented in Section 5 serves as a lower bound on program latency under resource constraints. The proofs for the following theorems are provided in Appendix A.